ratio of red marbles to blue is 3 to 7. If there are 250 marbles overall, how many are red? Need to show work. I have 3r/7b=x/250, but I don't think that's right.
Thank you!

Answer by ptaylor(2052)

Ok
If the ratio of red marbles to blue marbles is 3 to 7, that means that out of every 7+3=10 marbles 3 are red and 7 are blue. In other words, 3/10 are red and 7/10 are blue, soooo if there are 250 marbles, 3/10 are red and 7/10 are blue.
3/10(250)=75 are red and
7/10(250)=175 are blue and the ratio of red to blue is 75/175=3/7
Hope this helps--ptaylor